# On the Generalization of SFT: A Reinforcement Learning Perspective with Reward Rectification

### v1

@@ Abstract
We present a simple yet theoretically motivated improvement to Supervised Fine-Tuning (SFT) for the Large Language Model (LLM), addressing its limited generalization compared to reinforcement learning (RL). Through mathematical analysis, we reveal that standard SFT gradients implicitly encode a problematic reward structure that may severely restrict the generalization capabilities of model. To rectify this, we propose Dynamic Fine-Tuning (DFT), stabilizing gradient updates for each token by dynamically rescaling the objective function with the probability of this token. Remarkably, this single-line code change significantly outperforms standard SFT across multiple challenging benchmarks and base models, demonstrating greatly improved generalization. Additionally, our approach shows competitive results in offline RL settings, offering an effective yet simpler alternative. This work bridges theoretical insight and practical solutions, substantially advancing SFT performance. The code will be available at https://github.com/yongliang-wu/DFT.
@@ 1 Introduction
Supervised Fine-Tuning (SFT), which involves training a model on expert demonstration datasets, has become a standard Large Language Model (LLM) post-training approach for adapting models to new tasks or enhancing existing capabilities (Chung et al., 2024; Zhang et al., 2024; Sanh et al., 2022; Ouyang et al., 2022). Its widespread adoption is largely due to its ease of implementation and rapid acquisition of expert-like behaviors (Wei et al., 2022; Zhou et al., 2023). However, despite these advantages, SFT typically suffers from limited generalization compared to reinforcement learning (RL) methods (Chu et al., ; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Huan et al., 2025; Swamy et al., 2025). RL leverages explicit reward or verification signals, allowing the model to explore diverse strategies, thereby achieving stronger generalization. Nonetheless, RL approaches often require significant computational resources, are sensitive to hyperparameter tuning, and rely on the availability of reward signals, conditions that are not always feasible in practice (Schulman et al., 2017; Ouyang et al., 2022; Sheng et al., 2025; Strubell et al., 2019; Liu & Yin, 2024; Winsta, 2025). Even when RL is viable, SFT remains advantageous in rapidly acquiring expert behavioral patterns that RL may struggle to discover independently (Mandlekar et al., 2022; Chen et al., 2025b).
@@
To leverage the complementary strengths of both approaches, numerous hybrid methods have been developed by incorporating SFT and RL (Ouyang et al., 2022; Sheng et al., 2025; Rafailov et al., 2023; Liu et al., 2025; Qiu et al., 2025). However, can we fundamentally improve SFT itself? This is crucial since SFT is the only viable approach when there are no negative samples in the dataset and no reward or verification model available.
@@
In this work, we address this gap by providing a mathematical analysis that elucidates the fundamental differences between SFT and RL. We demonstrate that the gradient update of SFT can be interpreted as a special case of a policy gradient method with a specific, implicitly defined reward structure. Our analysis reveals that this implicit reward is both extremely sparse and inversely proportional to the policy’s assigned probability of expert actions (see Equation 6 for precise math details). Consequently, this leads to an ill-posed reward structure, especially when the model assigns low probabilities to expert actions; the resulting gradient experiences unbounded variance, creating a pathological optimization landscape.
@@
Building on this mathematical analysis, we propose Dynamic Fine-Tuning (DFT), a principled solution that addresses the root cause of the implicitly ill-posed reward structure. For each token, our method simply rescales the standard SFT objectives with the token probability, effectively neutralizing the inverse probability weighting that leads to an unexpected reward structure and unbounded variance. This theoretically motivated modification fundamentally transforms the gradient estimator from an unstable, biased and probability-dependent mechanism into a stable, uniformly-weighted update procedure.
The empirical results of our method are highly significant. For example, by fine-tuning the Qwen-2.5-Math models (Qwen Team et al., 2024b) on the NuminaMath dataset (LI et al., 2024), the improvement with our approach over the baseline is generally multiple times greater than that of improvement by SFT. Crucially, when standard SFT experiences performance degradation on challenging datasets including Olympiad Bench(International Mathematical Olympiad, 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023), likely due to overfitting, our method consistently deliver substantial improvements, highlighting its more robust generalization capabilities. The result is validated across model types, sizes, and data sizes (Table 1, Figure 1).
In addition, we explore the applicability of our method within RL scenarios (see Table 3, where negative samples or dense reward signals are available (Levine et al., 2020). Our experiments demonstrate that our approach not only significantly outperforms other offline RL methods such as DPO (Rafailov et al., 2023), RFT/RAFT (Dong et al., 2023; Ahn et al., 2024), but also our method performs favorably against online RL methods such as GRPO and PPO for Qwen2.5-Math-1.5B model in math tasks. Note that our method is much simpler in computation resources or procedure than most of the RL methods that may require a reference model or certain batch size, offering a practical alternative.
To understand how DFT affects the model differently, we analyze the change in the probability distribution after training (Figure  2. We observe that traditional SFT training uniformly increases token probabilities toward fitting the training data more tightly, whereas our approach, interestingly, also pushes some token distribution away from the training set. In particular, although most tokens still fit the training set closer, the percentage of less strongly fitted tokens increases markedly. We provide an in-depth discussion of this phenomenon in Section 4.3.1.
@@
The contributions of this work are theoretical and practical. On the theoretical side, we mathematically establish LLM SFT as a special RL in policy gradient space, pinpoint the underlying reasons for SFT’s limited generalization, and derive a method to improve it. On the experimental side, we show that such a simple solution, just one line of code, can substantially enhance LLM SFT’s performance and generalization capabilities across various tasks and models.
@@ 2 Related Work
The trade-off between SFT and RL is a central theme in modern language model alignment. SFT is widely adopted for its simplicity and efficiency in mimicking expert demonstrations (Chung et al., 2024; Zhou et al., 2023; Wei et al., 2022), a process analogous to behavioral cloning in robotics (Sammut, 2011; Mandlekar et al., 2022). However, the literature frequently notes that this approach can lead to overfitting and poor generalization compared to RL, which leverages reward signals to explore and discover more robust policies (Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Swamy et al., 2025). Recently, Chu et al. conduct a systematic comparison of SFT and RL on both textual and visual tasks, confirming that ”SFT memorizes while RL generalizes”. More importantly, they also find that SFT remains necessary as an initialization step to stabilize output formatting before RL training can be effective. Nevertheless, RL comes with significant practical hurdles, including high computational costs, hyperparameter sensitivity, and the need for an explicit reward function, often limit its applicability (Schulman et al., 2017; Strubell et al., 2019; Sheng et al., 2025).
@@
To exploit the benefits of both paradigms, a dominant line of research has focused on hybrid methods. The most established strategy involves an SFT pre-training phase followed by RL-based refinement, often using a learned reward model, as popularized by InstructGPT (Ouyang et al., 2022). More recent approaches have explored alternative combinations, such as interleaving SFT and RL steps to improve stability and performance (Sheng et al., 2025; Liu et al., 2025; Qiu et al., 2025). Other prominent approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), bypass explicit reward modeling by directly optimizing a policy on preference data, effectively integrating imitation and reinforcement signals into a single loss function. Chen et al. (2025a) Negative-aware Fine-Tuning (NFT) while enables LLMs to self-improve by modeling their own incorrect generations through an implicit negative policy. Although powerful, these methods are designed for settings where a reward signal, preference pairs, or negative samples are available. They augment the training pipeline, but do not fundamentally improve the SFT process in its native context, where there are only positive expert demonstrations. Our work diverges by focusing on enhancing SFT itself, without requiring any external feedback.
@@
On the other side, a theoretical line of inquiry has sought to unify SFT and RL. For example, Du et al. (2025) reframe RLHF as a reward-weighted form of SFT, simplifying the pipeline while maintaining dependence on an explicit reward. Wang et al. (2025) demonstrate that SFT can be viewed as an RL method with an implicit reward, proposing solutions such as smaller learning rates to manage an otherwise vanishing KL constraint. Abdolmaleki et al. (2025) analyze learning from both positive and negative feedback, showing how their balance affects policy convergence. Qin & Springenberg (2025) reframe SFT as a lower bound of RL and improve it by introducing importance weighting based on the data-generating policy. Although these works point out the general connection between SFT and RL through the lens of weighting, they stop short of providing a precise mathematical equivalence between the SFT gradient and the off-line policy gradient. In contrast, our work is the first to rigorously establish this equivalence, explicitly identifying that the key difference lies in an inverse-probability weighting term present in SFT. This insight directly motivates our proposed solution: simply multiplying the SFT loss with the model probability to cancel out the weighting.
Interestingly, our method yields a design for the cross-entropy (CE) loss that is diametrically opposite to the well-known Focal Loss (Lin et al., 2017). Our modified CE is $p\log(p)$, while the focal loss is $(1-p)^{\gamma}\log(p)$. Focal Loss intentionally downweights well-classified samples to improve performance on underrepresented classes, whereas we intentionally downweight poorly-classified samples to improve generalization. This contrast may reflect a fundamental shift in the era of LLMs, where underfitting becomes less problematic than overfitting.
@@ 3.1 Preliminaries
Let $\mathcal{D}=\{(x,y^{\star})\}$ denote a corpus of expert demonstrations, where $y^{\star}$ is the full reference response for query $x$. SFT minimizes the sentence-level cross-entropy:
@@
3.2 Unify SFT–RL Gradient Expression
@@
Define the auxiliary variables
$w(y\mid x)=\frac{\mathbf{1}}{\pi_{\theta}(y\mid x)},\quad r(x,y)=\mathbf{1}[y=y^{\star}],$
Reorganize the Equation 5 and rewrite it using the above auxiliary variables, we obtain the form
@@
This form of SFT gradient now closely aligns with policy gradient Equation 4. Thus we can see, conventional SFT is precisely an on-policy-gradient with the reward as an indicator function of matching the expert trajectory but biased by an importance weighting $1/\pi_{\theta}$.
Given the unavoidable sparsity of reward signals in the SFT setting, we identify the importance sampling weight $1/\pi_{\theta}$ as a fundamental cause of SFT’s poor generalization relative to RL. When the model assigns low probability to the expert response, the weight $w$ becomes large, resulting in unbounded and high-variance reward estimates from an RL perspective. This large variance issue is exacerbated by the extreme sparsity of the reward function—since $r(x,y)=\mathbf{1}[y=y^{\star}]$ is nonzero only when the model exactly matches the expert output. As a result, optimization tends to overfit to rare exact-match demonstrations, undermining the model’s ability to generalize beyond the training data.
@@ 3.3 Proposed Method
To neutralize the skewed reward issue identified when viewing SFT under an RL objective, we dynamically reweight the reward by multiplying by a corrective inverse ratio given by the policy probability $1/w$. The resulting “dynamically fine-tuned” gradient is then
@@
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator, ensuring that gradients do not flow through the reward scaling term $w$. To facilitate transitioning to later equations, we directly write $1/w$ to be $\pi_{\theta}(y^{\star}\mid x)$ instead of $\pi_{\theta}(y\mid x)$ because the indicator function in Equation 5 or Equation 6 would leave all cases where $y\neq y^{\star}$ as 0. Now since the gradient does not flow, the corrected SFT loss also becomes a simple reweighted loss, called Dynamic Fine-tuning (DFT).
$\mathcal{L}_{\text{DFT}}(\theta)=\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\Bigl{[}\operatorname{sg}\big{(}\pi_{\theta}(y^{\star}_{t}\mid x)\big{)}\log\pi_{\theta}(y^{\star}_{t}\mid x)\Bigr{]}.$ (8)
In practice, however, computing importance weights over the entire trajectory can induce numerical instability. A common treatment of this issue is to simply apply importance sampling in token-level, as was adopted in PPO (Schulman et al., 2017). This leads to the final DFT loss version:
@@ 4 Experiments
4.1 Main Experiment - SFT Setting
We focs on the standard SFT setting, characterized by having only expert demonstration data without negative samples, reward models, or verification signals. Our expert dataset typically originates from external policies, such as expert models or human annotations. The primary objective of this experiment is to rigorously evaluate whether DFT can robustly surpass standard SFT across diverse tasks, model architectures, model sizes, and dataset sizes.
@@ 4.1.1 Setup and Implementation Details
Dataset and Models.
We train with the NuminaMath CoT dataset (LI et al., 2024), comprising around 860,000 mathematical problems paired with the corresponding solutions. The dataset spans various sources, including Chinese high school mathematics exercises and U.S. and international mathematical olympiads. To efficiently manage computational resources, we randomly sample 100,000 instances from the dataset for training, which is sufficient since the evaluation accruacy curve 1 indicates the convergence of all methods well before the dataset exhaustion. We conduct experiments using multiple state-of-the-art models, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B (Qwen Team et al., 2024a), LLaMA-3.2-3B, LLaMA-3.1-8B (Dubey et al., 2024), and DeepSeekMath-7B-Base (Shao et al., 2024).
Training Details.
Our implementation builds upon the verl framework (Sheng et al., 2025), using recommended SFT hyperparameters. Specifically, we employ the AdamW optimizer with learning rates of $5\times 10^{-5}$ for all models except the LLaMA-3.1-8B, for which we adopt a lower learning rate of $2\times 10^{-5}$. We set the mini-batch size to 256 and the maximum input length to 2048 tokens. The learning rate follows a cosine decay schedule with a warm-up ratio of 0.1. We also include a concurrent method, Importance-Weighted SFT (iw-SFT) (Qin & Springenberg, 2025), for comparison. All training settings follow those reported in the original paper, except that we set the number of training epochs to 1.
Evaluation Settings.
For mathematical reasoning tasks, we evaluate on established benchmarks including Math500 (Hendrycks et al., ), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (AI Mathematical Olympiad, 2024), AIME 2024(American Institute of Mathematics, 2024), and AMC 2023(Mathematical Association of America, 2023). Each model uses the default chat template and Chain-of-Thought (CoT) prompting to stimulate step-by-step reasoning. All reported results represent average accuracy across 16 decoding runs, evaluated with a temperature of 1.0 and maximum generation length of 4096 tokens.
Table 1: Average@16 accuracy of five state-of-the-art large language models on five mathematical reasoning benchmarks: Math500, Minerva Math, Olympiad Bench, AIME 2024, and AMC 2023. The best performance of each model across benchmarks is bold.

### v2

@@ Abstract
In this work, we present a simple yet theoretically motivated improvement to Supervised Fine-Tuning (SFT) for the Large Language Model (LLM), addressing its limited generalization compared to reinforcement learning (RL). Through mathematical analysis, we reveal that standard SFT gradients implicitly encode a problematic reward structure that may severely restrict the generalization capabilities of model. To rectify this, we propose Dynamic Fine-Tuning (DFT), stabilizing gradient updates for each token by dynamically rescaling the objective function with the probability of this token. With just a single-line change, the method outperforms standard SFT on multiple difficult benchmarks and base models, from math reasoning to code generation and multi-modal tasks, demonstrating improved generalization. Additionally, DFT achieves competitive results in offline RL settings, providing an effective yet streamlined alternative. By bridging theoretical insights with practical solutions, this work advances the state of SFT. The code will be available at https://github.com/yongliang-wu/DFT.
@@ 1 Introduction
Supervised Fine-Tuning (SFT), which adapts models to expert demonstrations, has become the standard post-training paradigm for Large Language Models (LLMs). It enables efficient task adaptation and capability enhancement (Chung et al., 2024; Zhang et al., 2024b; Sanh et al., 2022; Ouyang et al., 2022), and is popular for its ease of implementation and rapid acquisition of expert-like behaviors (Wei et al., 2022; Zhou et al., 2023). Despite these advantages, SFT often shows limited generalization compared to reinforcement learning (RL) (Chu et al., 2024; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Huan et al., 2025; Swamy et al., 2025). RL leverages explicit reward or verification signals to explore diverse strategies and thus generalizes better. However, RL requires substantial computation, careful hyperparameter tuning, and explicit reward signals—conditions often impractical in real-world settings (Schulman et al., 2017; Ouyang et al., 2022; Sheng et al., 2025; Strubell et al., 2019; Liu & Yin, 2024; Winsta, 2025). Moreover, RL can struggle to recover expert-like behaviors that SFT captures efficiently (Mandlekar et al., 2022; Chen et al., 2025b).
@@
To exploit the complementary strengths of both approaches, many hybrid methods combine SFT with RL (Ouyang et al., 2022; Sheng et al., 2025; Rafailov et al., 2023; Liu et al., 2025; Qiu et al., 2025). Yet a key question remains: can SFT itself be fundamentally improved? This is crucial, as SFT remains the only viable option when datasets contain only positive demonstrations, with no negative samples or reward model available.
@@
In this work, we address this gap with a mathematical analysis of the connection between SFT and RL. We show that SFT’s gradient update is equivalent to a policy gradient with a specific, implicitly defined reward. Crucially, this reward is (i) sparse, and (ii) inversely proportional to the model’s probability of expert actions (see equation 6). As a result, when the model assigns low probability to expert actions, the gradient becomes excessively large—yielding an ill-posed reward structure and unstable, overfitting-prone optimization.
@@
Building on this insight, we propose Dynamic Fine-Tuning (DFT), a principled fix. Our method rescales the SFT objective at each token by its probability, canceling the distortion introduced by inverse-probability weighting. This transforms the SFT gradient from an unstable, biased estimator into a stable, uniformly weighted update rule.
Empirically, DFT delivers substantial improvements. On the Qwen-2.5-Math series (Qwen Team et al., 2024b) fine-tuned with NuminaMath-CoT (LI et al., 2024), DFT yields gains several times larger than standard SFT. More importantly, unlike SFT, which often degrades on challenging benchmarks such as OlympiadBench (He et al., 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023), our method consistently improves performance and generalization. These improvements hold across models, scales, and data sizes (Table 1, Figure 1), and extend to code generation and multimodal reasoning (Tables 3, 4).
We further test DFT in off-policy RL settings (Table 2), where dense rewards are available (Levine et al., 2020). Our method not only outperforms offline RL approaches such as DPO (Rafailov et al., 2023) and RAFT (Dong et al., 2023; Ahn et al., 2024), but also achieves competitive or superior performance to online methods like GRPO and PPO on math tasks with Qwen2.5-Math-1.5B. Unlike these RL methods, DFT requires neither a reference model nor large batch sizes, making it a simpler and more resource-efficient alternative.
To understand its effect, we analyze token probability distributions after training (Figure 2). While traditional SFT uniformly pushes probabilities toward the training set, DFT selectively increases some while reducing others. In particular, the proportion of less strongly fitted tokens rises, suggesting improved regularization and reduced overfitting. We provide further discussion in Appendix A.4.
@@
The contributions of this work are theoretical and practical. On the theoretical side, we mathematically establish LLM SFT as a special RL in policy gradient space, pinpoint the underlying reasons for the limited generalization of SFT, and derive a method to improve it. On the experimental side, we show that such a simple solution, just one line of code, can enhance the performance and generalization capabilities of SFT across various tasks and models.
@@ 2 Related Work
The trade-off between supervised fine-tuning (SFT) and reinforcement learning (RL) is central to the alignment of large language models. SFT is widely adopted due to its simplicity and efficiency in imitating expert demonstrations (Chung et al., 2024; Zhou et al., 2023; Wei et al., 2022), analogous to behavioral cloning in robotics (Sammut, 2011; Mandlekar et al., 2022). However, the literature consistently highlights its limitations, particularly the tendency to overfit and generalize poorly compared to RL, which leverages reward signals to discover more robust policies (Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Swamy et al., 2025; Zhang et al., 2025). A recent systematic comparison by Chu et al. (2024) across textual and visual domains confirms this distinction, concisely summarized as “SFT memorizes while RL generalizes.” They further show that SFT remains indispensable as an initialization step, stabilizing output formatting prior to effective RL training. Nonetheless, RL faces significant practical hurdles, including computational expense, sensitivity to hyperparameters, and the requirement of an explicit reward function, all of which constrain its applicability (Schulman et al., 2017; Strubell et al., 2019; Sheng et al., 2025).
@@
To combine the strengths of both paradigms, much recent work has pursued hybrid approaches. The most common strategy involves SFT pretraining followed by RL-based refinement with a learned reward model, as popularized by InstructGPT (Ouyang et al., 2022). More recent methods interleave SFT and RL updates to improve stability and performance (Sheng et al., 2025; Liu et al., 2025; Qiu et al., 2025). Other approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), bypass reward modeling entirely by directly optimizing policies on preference data, thereby unifying imitation and reinforcement signals within a single loss function. Chen et al. (2025a) introduce Negative-aware Fine-Tuning (NFT), which models incorrect generations via an implicit negative policy, enabling self-improvement without explicit feedback. While powerful, these methods rely on reward signals, preference pairs, or negative samples. They enrich the training pipeline but do not fundamentally improve SFT in its native setting, where only positive demonstrations are available. Our work instead focuses on enhancing SFT itself without requiring external feedback.
@@
A complementary line of theoretical research seeks to unify SFT and RL under a common formalism. Du et al. (2025) reinterpret RLHF as a reward-weighted variant of SFT, preserving reliance on an explicit reward. Wang et al. (2025) show that SFT can be cast as RL with an implicit reward, proposing adjustments such as smaller learning rates to manage the vanishing KL constraint. Abdolmaleki et al. (2025) analyze learning from both positive and negative feedback, studying how their balance affects convergence. Qin & Springenberg (2025) view SFT as a lower bound of RL and introduce importance weighting based on the data-generating policy. While these works establish connections between SFT and RL through weighting, they do not provide a precise mathematical equivalence between the SFT gradient and the offline policy gradient. Some methods approximate this connection in practice by reweighting training losses. For instance, MixCE (Zhang et al., 2023) combines the forward and reverse KL divergences to form a unified objective, while GOLD (Pang & He, 2021) adopts offline RL with demonstrations, introducing reliance on an unknown demonstration distribution $\pi_{b}$ and a restrictive $1/N$ assumption. In contrast, our work rigorously establishes this equivalence, pinpointing the inverse-probability weighting term in SFT as the key distinction. This observation directly motivates our proposed solution: multiplying the SFT loss by the model probability to eliminate the weighting.
Interestingly, our method modifies the standard cross-entropy (CE) loss in a way that inverts the weighting philosophy of the widely used Focal Loss (Lin et al., 2017). Specifically, our modified CE takes the form $p\log(p)$, whereas focal loss is defined as $(1-p)^{\gamma}\log(p)$. Focal Loss deliberately downweights well-classified samples to emphasize underrepresented or hard cases, whereas we deliberately downweight poorly classified samples to encourage generalization. This inversion reflects a fundamental shift in the LLM era: while underfitting was once a central challenge, overfitting and memorization now dominate, demanding a rethinking of objective design.
@@ 3.1 Preliminaries
Let $\mathcal{D}=\{(x,y^{\star})\}$ denote a corpus of expert demonstrations, where $y^{\star}$ is the complete reference response to the query $x$. SFT minimizes the sentence-level cross-entropy:
@@
3.2 Unify SFT and RL Gradient Expression
@@
Define the auxiliary variables (importance sampling weight) as
$w(y\mid x)=\frac{\mathbf{1}}{\pi_{\theta}(y\mid x)},\quad r(x,y)=\mathbf{1}[y=y^{\star}].$
Reorganizing equation 5 and rewriting it using the above auxiliary variables, we obtain the form
@@
This form of the SFT gradient now aligns closely with the policy gradient equation 4. Thus we can see, conventional SFT is precisely an on-policy-gradient with the reward as an indicator function of matching the expert trajectory but biased by an importance weighting $1/\pi_{\theta}$. More detailed derivations can be found in Appendix A.3.
Given the unavoidable sparsity of reward signals in the SFT setting, we identify the importance sampling weight $1/\pi_{\theta}$ as a fundamental cause of the poor generalization of SFT relative to RL. When the model assigns low probability to the expert response, the weight $w$ becomes excessively large, leading to an ill-posed reward structure. In such cases, the gradient can grow disproportionately, making the training unstable and prone to overfitting. This issue is further exacerbated by the extreme sparsity of the reward function, since $r(x,y)=\mathbf{1}[y=y^{\star}]$ is non-zero only when the model exactly matches the expert output. As a result, optimization tends to overfit rare exact-match demonstrations, undermining the model’s ability to generalize beyond training data.
@@ 3.3 Proposed Method
To neutralize the skewed reward issue identified when viewing SFT under the RL objective, we dynamically reweight the reward by multiplying by a corrective inverse ratio given by the policy probability $1/w$. The resulting “dynamically fine-tuned” gradient is then
@@
where $\operatorname{sg}(\cdot)$ denotes the stop gradient operator, ensuring that gradients do not flow through the reward scaling term $w$. To facilitate transition to later equations, we directly write $1/w$ to be $\pi_{\theta}(y^{\star}\mid x)$ instead of $\pi_{\theta}(y\mid x)$ because the indicator function in equation 5 or equation 6 would leave all cases where $y\neq y^{\star}$ is 0. Now since the gradient does not flow, the corrected SFT loss also becomes a simple reweighted loss, called Dynamic Fine-tuning (DFT).
$\mathcal{L}_{\text{DFT}}(\theta)=\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\Bigl[\operatorname{sg}\big(\pi_{\theta}(y^{\star}\mid x)\big)\log\pi_{\theta}(y^{\star}\mid x)\Bigr].$ (8)
However, in practice, computing importance weights over the entire trajectory can induce numerical instability. A common treatment of this issue is to simply apply importance sampling at the token level, as was adopted in PPO (Schulman et al., 2017). This leads to the final DFT loss version:
@@ 4 Experiments
We design four groups of experiments to comprehensively evaluate DFT. We first study the standard SFT setting on mathematical reasoning tasks to establish its core advantage over SFT (Section 4.1). We then extend to an offline RL setting, comparing DFT with representative offline and online RL methods (Section 4.2). To test cross-domain robustness, we further examine DFT on code generation benchmarks (Section 4.3) and its applicability to multi-modal reasoning math datasets (Section 4.4).
4.1 Main Experiment - Mathematical Reasoning Task
To examine whether DFT can outperform vanilla SFT across tasks, architectures, and scales, we use mathematical reasoning as a representative testbed.
@@ 4.1.1 Setup and Implementation Details
To efficiently manage computational resources, We andomly sample 100,000 instances from the the NuminaMath-CoT dataset (LI et al., 2024) for training. We conduct experiments using multiple models, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B (Qwen Team et al., 2024a), LLaMA-3.2-3B, LLaMA-3.1-8B (Dubey et al., 2024), and DeepSeekMath-7B (Shao et al., 2024).
Our implementation builds upon the verl framework (Sheng et al., 2025), using recommended SFT hyper-parameters. Specifically, we employ the AdamW optimizer with learning rates of $5\times 10^{-5}$ for all models except the LLaMA-3.1-8B-Base, for which we adopt a lower learning rate of $2\times 10^{-5}$. We set the mini-batch size to 256 and the maximum input length to 2048 tokens. The learning rate follows a cosine decay schedule with a warm-up ratio of 0.1. We evaluate on benchmarks including Math500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (He et al., 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023) through the official Qwen2.5-Math evaluation pipeline (Qwen Team et al., 2024a). Each model uses the default chat template and Chain-of-Thought (CoT) prompting to stimulate step-by-step reasoning. All reported results represent average accuracy across 16 decoding runs, evaluated with a temperature of 1.0 and maximum generation length of 4096 tokens.
Table 1: Average@16 accuracy of five state-of-the-art large language models on mathematical reasoning benchmarks. The best performance of each model across benchmarks is bold.